# CAT-SAM: Conditional Tuning Network for Few-Shot Adaptation of Segmentation Anything Model

### v1

@@ 3.3 Training and Inference
During training, we feed target samples together with geometric prompts into CAT-SAM to generate CAT-SAM masks, under the supervision of ground-truth masks as well as a linear combination of BCE loss and dice loss [29]. We freeze the pre-trained SAM parameters and update solely the parameters that are introduced in the tuning modules described above. Note we train CAT-SAM with a mixture of sampled geometric prompts, including bounding boxes, randomly selected points, and coarse masks. In addition, we create degraded coarse masks by introducing random Gaussian noise in the boundary regions of the ground-truth masks as in [17]. Further training details can be found in Section 4.1 and the appendix.
During the inference phase, the input image is fed into SAM’s image encoder along with the integrated prompt tokens in CAT-SAM-T or adapters in CAT-SAM-A, to generate adaptive image embeddings. These embeddings, combined with the prompt tokens from SAM’s prompt encoder, serve as input to the mask decoder. Subsequently, the HQ-Output token and associated MLP are utilized for the target mask prediction. Lastly, we up-sample the mask to the original resolution to produce the final output.
4 Experiments
We evaluate CAT-SAM’s efficacy through comprehensive experiments involving eight segmentation tasks across 11 datasets, all hailing from challenging downstream fields that SAM struggles to address effectively. Our experiments encompass a wide array of scenarios and tasks, spanning from varying target shots (from one-shot to full-shot adaptation) to different imagery modalities.
4.1 Experimental Setup
Datasets. We conduct experiments on two collections of datasets listed in Table 2: 1) True-color images, including WHU [13] for building segmentation, Kvasir [32] for polyp segmentation, SBU-Shadow [42] for shadow segmentation, Massachusetts (MA.) Roads [30] for road segmentation, as well as fine-grained segmentation datasets DIS [33], ThinObject [26], HRSOD [47], and COIFT [26]. 2) Non-RGB domains, encompassing JSRT [40] for chest organ segmentation (X-ray images), HRSID [44] for ship instance segmentation (SAR images), and FLS [41] for marine debris segmentation (Sonar images). We use official dataset splits for fair comparisons.
Evaluation Metrics. For single-class foreground datasets including WHU, Kvasir, SBU-Shadow, and MA. Roads, we employ standard mask IoU. JSRT and FLS, which have multiple classes, are evaluated using individual class IoU and their average. As for DIS, ThinObject, HRSOD, and COIFT, we follow [17] and report mIoU and the boundary metrics (mBIoU) [6] for fair comparisons. For instance segmentation of HRSID, we use standard AP, AP50, and AP75.
@@ 4.1 Experimental Setup
Table 3: Ablation study of CAT-SAM-T in (a) and CAT-SAM-A in (b) for 1-shot adaptation over datasets WHU [13] (on building segmentation), Kvasir [32] (on polyp segmentation), and SBU-Shadow [42] (on shadow segmentation). “Enc-T.” and “Enc-A.” means tuning the image encoder with prompt tokens and adapters, respectively. “Dec.” means tuning the mask decoder. “PB-T” and “PB-A” denote two customized prompt bridges for conditional tuning. Model 1 is the baseline SAM without adaptation.
Implementation Details. Unless otherwise specified, CAT-SAM and its comparing models use ViT-L [9] as the image encoder backbone. We train on 1 NVIDIA RTX A6000 GPU for one-shot adaptation, while 4 GPUs for multiple-shots. Following [17], ground truth boxes serve as the default input geometric prompts during evaluation to ensure fair comparison and minimize randomness, with the exception of Figure 3 where point prompts are evaluated. Please refer to the appendix for more implementation details.
4.2 Ablation Studies
We first investigate the contributions of different tuning modules within both CAT-SAM variants to assess their impact on overall adaptation performance. Table 3 shows the experiments of one-shot adaptation across three true-color image segmentation benchmarks WHU [13], Kvasir [32], and SBU-Shadow [42]. For each CAM-SAM variant CAM-SAM-T/CAM-SAM-A, we compare 5 models including: 1. The original SAM without adaptation (baseline); 2. Tuning SAM’s image encoder alone with prompt tokens or adapters; 3. Tuning SAM’s mask encoder alone; 4. Tuning SAM’s image encoder and mask decoder independently; and 5. Conditional tuning of SAM’s image encoder and mask decoder with the proposed prompt bridging (i.e., the complete CAT-SAM-T and CAT-SAM-A).
@@ 4.2 Ablation Studies
Methods WHU Kvasir Shadow AVG
@@
Table 4: One-shot adaptation over challenging true-color datasets WHU [13] (on building segmentation), Kvasir [32] (on Polyp segmentation), and SBU-Shadow [42] (on shadow segmentation). The baseline is SAM without adaptation.
@@
CAT-SAM-T (ours) 84.0 78.1 95.6 92.0 93.4 87.6 94.0 87.9 91.7 86.4
CAT-SAM-A (ours) 83.6 77.7 95.6 92.2 93.1 87.1 94.1 88.2 91.6 86.3
4.3 Comparison with the State-of-the-Art
Since few-shot adaptation for SAM remains a relatively under-explored task with very limited prior studies, we evaluate our approach alongside several highly related adaptation methods including HQ-SAM [17], VPT [16], and AdaptFormer [3] in computer vision, as well as LoRA [12] for adapting large language models.
Table 4 shows experiments on one-shot adaptation of WHU, Kvasir, and SBU-Shadow. We can observe that AdaptFormer and LoRA exhibit notably better performance than VPT-deep and VPT-shallow, underscoring that modifying the model’s structure is a more effective approach than prompt tuning in SAM adaptation. It is noteworthy that all four adaptation methods outperform HQ-SAM, demonstrating the greater importance of tuning the image encoder compared to tuning the mask decoder. Such results can be largely explained by the fact that HQ-SAM is tailored for intricate natural images, for which SAM’s image encoder is well-trained. As a comparison, we focus on more challenging unnatural images, where SAM’s image encoder faces difficulties, necessitating image encoder tuning for adaptive feature representations. The two proposed CAT-SAM variants consistently outperform all compared methods, with CAT-SAM-A edging slightly due to the modification of network structures. This verifies the superiority of CAT-SAM in SAM adaptation with few-shot target samples.
@@ 4.3 Comparison with the State-of-the-Art
We also examine CAT-SAM in the task of intricate structural image segmentation. To ensure fair comparisons, we follow [17] and train CAT-SAM variants on the full HQSeg-44K dataset, which consists of 44,320 meticulously annotated image masks. Table 5 shows experiments across four exceptionally fine-grained segmentation datasets including DIS [33], ThinObject [26], HRSOD [47], and COIFT [26] (in mIoU and the boundary metrics (mBIoU) [6]). We can see that CAT-SAM-T achieves slightly superior performance compared to CAT-SAM-A, while both clearly outperform HQ-SAM and the original SAM. This serves as additional evidence of the efficacy of our proposed conditional tuning in adapting SAM toward challenging downstream tasks.
4.4 CAT-SAM with Varying Target Samples
We also investigate how varying numbers of target training samples affect the adaptation performance of the proposed CAT-SAM. Specifically, we randomly select either one target image or sixteen target images for 1-shot and 16-shot adaptations, respectively, while employing the entire training set for full-shot adaptation.
True-Color Datasets. Table 6 shows experiments on WHU, Kvasir, and SBU-Shadow. It can be observed that the inclusion of more training samples consistently enhances target segmentation across the three datasets. Notably, even with just one-shot target samples, the two CAT-SAM variants achieve high-quality target segmentation at 82.7% and 85.2% mIoU across the three datasets, respectively. This corresponds to 90% and 93% of the performance attained with a full training set, where both CAT-SAM variants achieve 92.1% mIoU. The experimental results demonstrate the remarkable efficacy of CAT-SAM in few-shot scenarios, enabling much-reduced target samples in adaptation and significantly expanding SAM’s applicability across downstream tasks.
#Samples Methods WHU Kvasir Shadow AVG
None SAM [19] 43.5 79.0 62.4 61.6
1-shot CAT-SAM-T 86.8 83.4 78.0 82.7
CAT-SAM-A 88.2 85.4 81.9 85.2
16- CAT-SAM-T 89.6 88.5 86.3 88.1
shot CAT-SAM-A 90.3 89.3 87.5 89.0
Full- CAT-SAM-T 93.3 90.8 92.3 92.1
shot CAT-SAM-A 93.6 90.5 92.3 92.1
Table 6: Adapting two CAT-SAM variants across challenging true-color image datasets with a varying number of training samples.
#Samples Methods JSRT [40] FLS [41]
Lungs Heart mIoU Bott. Can Chain Drin. Hook Prop. Sham. Stan. Tire Valve Wall mIoU
None SAM [19] 85.0 71.9 78.5 76.0 81.4 38.4 78.5 47.9 71.6 81.7 85.7 69.3 54.0 82.7 69.7
1-shot CAT-SAM-T 95.4 90.5 93.0
CAT-SAM-A 94.7 90.5 92.6
16-shot CAT-SAM-T 95.8 92.7 94.2 78.1 76.7 46.4 82.3 49.5 73.5 82.8 85.6 83.2 61.8 85.0 73.2
CAT-SAM-A 95.5 91.5 93.5 72.4 72.6 47.9 78.9 52.7 69.6 81.6 87.4 73.8 62.0 86.3 71.4
Full-shot CAT-SAM-T 96.3 92.6 94.4 86.3 87.9 66.7 83.2 67.0 79.5 87.0 88.7 90.9 70.3 91.3 81.7
CAT-SAM-A 96.4 92.8 94.6 85.5 88.1 67.1 83.5 67.6 80.2 86.5 89.8 91.2 69.9 92.1 82.0
Table 7: Few-shot adaptation with varying number of training samples for non-RGB datasets, including FLS [41] with Sonar images for marine debris segmentation and JSRT [40] for chest X-ray segmentation. 1-shot adaptation over FLS is unattainable due to the absence of images containing objects from all 11 classes.
Methods #Samples AP AP50 AP75
SAM [19] None 38.2 86.6 26.8
CAT-SAM-T 1-shot 46.0 88.4 45.4
CAT-SAM-A 44.9 88.0 41.4
CAT-SAM-T 16-shot 46.2 89.3 45.2
CAT-SAM-A 45.7 89.4 43.7
CAT-SAM-T Full-shot 51.4 93.0 55.0
CAT-SAM-A 52.9 93.9 56.0
Table 8: Few-shot adaptation with varying number of training samples on the HRSID [44] Dataset for Ship Instance Segmentation with SAR Images.
@@ 4.4 CAT-SAM with Varying Target Samples
Non-RGB Datasets. Table 7 shows experiments on JSRT [40] for chest organ segmentation with X-ray images, as well as FLS [41] for marine debris segmentation with Sonar images. The two datasets represent novel image distributions as compared with SAM’s training data. Note we train instances of all classes together while evaluating them by classes. Since no images in FLS contain objects of all 11 classes, we only report results of 16-shot and full-shot adaptation for FLS. As the table shows, CAT-SAM-T and CAT-SAM-A improve the mIoU to 93.0% and 92.6%, respectively, with only one shot of JSRT training image. They further achieve marginal gains of 1.4% and 2.0% with full-shot adaptation. These experimental results are well aligned with those reported in Table 6.
@@
Experiments on FLS [41] show certain different results. While both CAT-SAM variants exhibit clear adaptation effects, the performance gains for both 16-shot and full-shot adaptation are notably lower as compared with previously studied datasets. This discrepancy is largely attributed to two primary factors: 1) the inherent difficulty of FLS with its 11 semantic classes; and 2) the larger variation in object sizes compared to datasets evaluated previously. These experiments further underscore the need for a more comprehensive adaptation of SAM for exceptionally challenging downstream tasks. We expect further studies in this regard.
In Table 8, we also present the results of ship instance segmentation using the high-resolution SAR image dataset HRSID [44]. This dataset presents different challenges as ships are very small in size. The results show that CAT-SAM consistently outperforms SAM in this diverse domain, highlighting its efficiency in domain adaptation.

### v3

@@ 3.3 Training and Inference
During training, we feed target samples together with geometric prompts into CAT-SAM to generate CAT-SAM masks, under the supervision of ground-truth masks as well as a linear combination of BCE loss and dice loss [31]. We freeze the pre-trained SAM parameters and update solely the parameters that are introduced in the above-described tuning modules. Note we train CAT-SAM with a mixture of sampled geometric prompts, including bounding boxes, randomly selected points, and coarse masks. In addition, we create degraded coarse masks by introducing random Gaussian noise in the boundary regions of the ground-truth masks as in [21]. Further training details can be found in Section 4.1 and the appendix.
During the inference phase, the input image is fed into SAM’s image encoder along with the integrated prompt tokens in CAT-SAM-T or adapters in CAT-SAM-A, to generate adaptive image embeddings. These embeddings, combined with the prompt tokens from SAM’s prompt encoder, serve as input to the mask decoder. Subsequently, the updated CAT-token and associated MLP are utilized for the target mask prediction. Lastly, we up-sample the mask to the original resolution to produce the final output.
@@
4 Experiments
We evaluate CAT-SAM’s efficacy through comprehensive experiments involving eight segmentation tasks across 11 datasets, all hailing from challenging downstream fields that SAM struggles to address effectively. Our experiments encompass a wide array of scenarios and tasks, spanning from varying target shots to different imagery modalities.
4.1 Experimental Setup
Datasets. We conduct experiments on two collections of datasets listed in Tab. 2: 1) True-color images, including WHU [18] for building segmentation, Kvasir [34] for polyp segmentation, SBU-Shadow [42] for shadow segmentation, Massachusetts (MA.) Roads [32] for road segmentation, as well as fine-grained segmentation datasets DIS [35], ThinObject [28], HRSOD [51], and COIFT [28]. 2) Non-RGB domains, encompassing JSRT [39] for chest organ segmentation (X-ray images), HRSID [45] for ship instance segmentation (high-resolution SAR images), and FLS [40] for marine debris segmentation (Sonar images). We use official dataset splits for fair comparisons.
Evaluation Metrics. For single-class datasets including WHU, Kvasir, SBU-Shadow, and MA. Roads, we employ standard mask IoU of foreground. JSRT and FLS, which have multiple classes, are evaluated using individual class IoU and their average. As for DIS, ThinObject, HRSOD, and COIFT, we follow [21] and report mIoU and the boundary metrics (mBIoU) [7] for fair comparisons. For instance segmentation of HRSID, we use standard AP, AP50, and AP75.
Implementation Details. Unless otherwise specified, CAT-SAM and its comparing models use ViT-L [9] as the image encoder backbone. We train on 1 NVIDIA RTX A6000 GPU for one-shot adaptation, while 4 GPUs for multiple-shots. Following [21], ground truth boxes serve as the default input geometric prompts during evaluation to ensure fair comparison and minimize randomness, with the exception of Fig. 3 where point prompts are evaluated. Please refer to the appendix for more implementation details.
4.2 Ablation Studies
Table 3: Ablation study of CAT-SAM-T in (a) and CAT-SAM-A in (b) for 1-shot adaptation over datasets: "W" for WHU [18] (on building segmentation), "K" for Kvasir [34] (on polyp segmentation), and "S" for SBU-Shadow [42] (on shadow segmentation). "Enc-T." and "Enc-A." means tuning the image encoder with prompt tokens and adapters, respectively. "Dec." means tuning the mask decoder. "PB-T" and "PB-A" denote two customized prompt bridges for conditional tuning. Model 1 is the baseline SAM without adaptation.
@@ 4.2 Ablation Studies
We first investigate the contributions of different tuning modules within both CAT-SAM variants to assess their impact on overall adaptation performance. Tab. 3 shows the experiments of one-shot adaptation across three true-color image segmentation benchmarks WHU [18], Kvasir [34], and SBU-Shadow [42]. For each CAM-SAM variant CAM-SAM-T/CAM-SAM-A, we compare 5 models including: 1. The original SAM without adaptation (baseline); 2. Tuning SAM’s image encoder alone with prompt tokens or adapters; 3. Tuning SAM’s mask decoder alone; 4. Tuning SAM’s image encoder and mask decoder independently; and 5. Conditional tuning of SAM’s image encoder and mask decoder with the proposed prompt bridging (i.e., the complete CAT-SAM-T and CAT-SAM-A).
@@
4.3 Comparison with the State-of-the-Art
CAT-SAM for One-Shot Adaptation. To assess CAT-SAM’s effectiveness in reducing downstream training data and its adaptability with few-shot samples, we first evaluate it under the extremely challenging one-shot adaptation scenario. Given the lack of prior studies under such setup, we comprehensively benchmark it against closely related methods that can be broadly categorized into two groups. The first group comprises state-of-the-art tuning methods for foundation models of different modalities, including VPT [20] and AdaptFormer [3] for large vision foundation models, and LoRA [16] for large language models. The second group comprises recent SAM-based adaptation methods which were largely designed for full-shot adaptation [21, 46, 41, 53, 30] ([47] tackles few-shot adaptation but requires 20 shots).
Table 4: Comparison of adaptive segmentation performance on challenging true-color datasets: WHU [18] (on building segmentation), Kvasir [34] (on Polyp segmentation), and SBU-Shadow [42] (on shadow segmentation). The baseline is SAM without adaptation. All compared methods utilize one-shot sample for adaptation, except for those denoted by * which utilize 20 shots with metric numbers sourced from [47].
Methods WHU Kvasir SBU-Shadow Average
@@ 4.3 Comparison with the State-of-the-Art
AdaptFormer [3] 83.2 76.8 77.2 79.1
LoRA [16] 86.1 77.5 74.4 79.3
Med-SA [46] 34.5 28.6 22.1 28.4
MedSAM [30] 30.6 29.8 53.4 37.9
SimAda [41] 48.6 18.7 49.3 38.9
SPFS-SAM* [47] (20-shots) - 53.4 - -
SAMed* [53] (20-shots) - 51.7 - -
@@
Tab. 4 presents results over WHU, Kvasir, and SBU-Shadow datasets. We can observe that all three tuning methods for foundation models exhibit clear adaptation effects. Notably, AdaptFormer and LoRA demonstrate significantly better performance than VPT-deep and VPT-shallow, highlighting the effectiveness of modifying the model’s structure over prompt tuning in the input space during SAM adaptation. On the other hand, most SAM-based adaptation methods, including Med-SA, SimAda, SPFS-SAM, SAMed, and MedSAM, yield much lower performance, primarily due to their reliance on many more target images which leads to overfitting while learning from one-shot target sample. HQ-SAM achieves substantial segmentation improvements but it still cannot compete with the three tuning methods, demonstrating the significance of tuning the image encoder over the mask decoder. Finally, both CAT-SAM variants consistently outperform all compared methods, with CAT-SAM-A exhibiting a slight edge due to its network structure modifications. These experiments validate the superiority of the proposed decoder-conditioned joint tuning approach in extracting domain-specific features with few-shot samples.
Table 5: Comparison of state-of-the-art segmentation achieved by fully supervised learning (FSL) methods trained on full-shot data (upper part) and mask segmentation by SAM and our CAT-SAM (lower part). Evaluation follows official criteria of each FSL benchmark, reporting IoU for WHU, mIoU for Kvasir, and BER for SBU-Shadow to ensure consistency. "(1)", "(16)", and "(full)" after CAT-SAM-T/CAT-SAM-A denote tuning with 1-shot, 16-shots, and full-shots samples.
WHU Kvasir SBU-Shadow
Model IoU Model mIoU Model BER(↓)
STT [2] 90.5 CASCADE [36] 87.8 BDRAR [59] 3.64
BuildFormer [44] 91.4 PatchRefineNet [33] 89.1 DSDNet [56] 3.45
CBR-Net [14] 91.4 DUCK-Net [10] 90.5 MTMT [6] 3.15
SAM 43.5 SAM 87.3 SAM 13.22
CAT-SAM-T (1) 86.8 CAT-SAM-T (1) 90.0 CAT-SAM-T (1) 7.81
CAT-SAM-A (1) 88.2 CAT-SAM-A (1) 91.3 CAT-SAM-A (1) 5.27
CAT-SAM-T (16) 89.6 CAT-SAM-T (16) 93.1 CAT-SAM-T (16) 4.04
CAT-SAM-A (16) 90.3 CAT-SAM-A (16) 93.6 CAT-SAM-A (16) 3.80
CAT-SAM-T (full) 93.3 CAT-SAM-T (full) 94.5 CAT-SAM-T (full) 2.54
CAT-SAM-A (full) 93.6 CAT-SAM-A (full) 94.3 CAT-SAM-A (full) 2.39
Full-Shot Learning. While CAT-SAM is primarily designed for few-shot adaptation, it also demonstrates proficiency in full-shot adaptation scenarios. In Tab. 5, we present the state-of-the-art performances of traditional fully-supervised learning (FSL) methods trained over full-shot data and compare them with SAM and our CAT-SAM (adapted with various target samples). We report official evaluation metrics for three public benchmarks, including IoU for WHU, mIoU for Kvasir, and BER [17] for SBU-Shadow. Note the distinction in setups between the FSL segmentation and the mask segmentation: FSL focuses on traditional semantic segmentation of specific classes without prompt input while SAM and CAT-SAM involve general mask segmentation with prompt input. Despite these differences, these performances can serve as valuable references for evaluating CAT-SAM’s segmentation ability.
@@
CAT-SAM-T 84.0 78.1 95.6 92.0 93.4 87.6 94.0 87.9 91.7 86.4
CAT-SAM-A 83.6 77.7 95.6 92.2 93.1 87.1 94.1 88.2 91.6 86.3
With full target training data, both CAT-SAM variants outperform state-of-the-art FSL methods consistently. Notably, the two CAT-SAM variants achieve impressive target segmentation under the challenging one-shot setup with a single target sample. With 16 shots of target samples, their performance is even on par with the FSL methods that are trained by using full target training data. These experiments underscore the superiority of effective transfer of broad knowledge from the foundation model SAM as compared with training from scratch with FSL. It also demonstrates the remarkable efficacy of CAT-SAM in substantially expanding SAM’s applicability across various downstream tasks.
@@
We conducted further experiments to compare CAT-SAM against HQ-SAM [21], a state-of-the-art SAM-based adaptation solution for intricate structural image segmentation. For fair comparisons, we follow the evaluation protocol of HQ-SAM and tune CAT-SAM on the complete HQSeg-44K dataset, consisting of 44,320 meticulously annotated image masks. Tab. 6 shows experiments across four exceptionally fine-grained segmentation datasets including DIS [35], ThinObject [28], HRSOD [51], and COIFT [28] (in mIoU and the boundary metrics (mBIoU) [7]). We can see that CAT-SAM-T achieves slightly better performance than CAT-SAM-A, while both clearly outperform HQ-SAM and the original SAM. These experiments further demonstrate the efficacy of our proposed conditional tuning in adapting SAM toward challenging downstream tasks. Visual qualitative comparisons are provided in the appendix.
4.4 CAT-SAM for Non-RGB Domains
@@
We also explore CAT-SAM’s performance across more challenging image modalities characterized by larger domain discrepancies. Specifically, we assess its efficacy on non-RGB domains, including JSRT [39] for chest organ segmentation using X-ray images and FLS [40] for marine debris segmentation with Sonar images, and HRSID [45] for ship instance segmentation with SAR images. For adaptation, we randomly select either one or sixteen target images for 1-shot and 16-shot scenarios, respectively, while employing the entire training set for full-shot adaptation. These datasets represent novel image distributions as compared with SAM’s training data. Note we train instances of all classes together while evaluating them by classes. Since no images in FLS contain objects of all 11 classes, we only report results of 16-shot and full-shot adaptation for FLS.
Table 7: Few-shot adaptation with varying number of training samples for non-RGB datasets, including FLS [40] with Sonar images for marine debris segmentation, JSRT [39] for chest X-ray segmentation, and HRSID [45] for ship instance segmentation with SAR images. "N/A" denotes that 1-shot adaptation over FLS is unattainable due to the absence of images containing objects from all 11 classes.
#Samples Methods JSRT [39] FLS [40] HRSID [45]
Lungs Heart mIoU mIoU AP AP50 AP75
None SAM [23] 85.0 71.9 78.5 69.7 38.2 86.6 26.8
1-shot CAT-SAM-T 95.4 90.5 93.0 N/A 46.0 88.4 45.4
CAT-SAM-A 94.7 90.5 92.6 N/A 44.9 88.0 41.4
16-shot CAT-SAM-T 95.8 92.7 94.2 73.2 46.2 89.3 45.2
CAT-SAM-A 95.5 91.5 93.5 71.4 45.7 89.4 43.7
Full-shot CAT-SAM-T 96.3 92.6 94.4 81.7 51.4 93.0 55.0
CAT-SAM-A 96.4 92.8 94.6 82.0 52.9 93.9 56.0
@@
Tab. 7 shows experimental results. Notably, CAT-SAM-T and CAT-SAM-A improve the mIoU to 93.0% and 92.6%, respectively, with only one shot of JSRT training image. The improvements saturate gradually as the number of training samples increases, with marginal gains of 1.4% and 2.0%, respectively, when adapted with full-shot target data. These findings underscore the superior few-shot capability of the proposed CAT-SAM and align well with those reported in Tab. 5. Experiments on FLS [40] show different results. While both CAT-SAM variants exhibit clear adaptation effects, the performance gains for both 16-shot and full-shot adaptation are notably lower as compared with previously evaluated datasets. The discrepancy is largely attributed to two primary factors: 1) the inherent difficulty of FLS with its 11 semantic classes (per-class performances are provided in appendix); and 2) the larger variation of FLS data in shapes and object sizes as compared with previously evaluated data. As for ship instance segmentation using SAR images in HRSID [45], this dataset presents different challenges as ships are very small in size. The results show that CAT-SAM consistently outperforms SAM in this diverse domain, highlighting its efficiency in domain adaptation with large domain discrepancy. These experiments further underscore the need for a more comprehensive adaptation of SAM for exceptionally challenging downstream tasks. We expect further studies along this research direction.